Suppose the spot rate for the Euro is at $1.1327 per euro. What is the minimum value of an American Put option contract with a strike price of $1.1541 per euro? Assume a contract size of EUR15,000
Answer #1

The minimum value of a put option is the intrinsic value of the option.

For a put option, intrinsic value = Strike price - Spot price (when Spot price < strike price)

Intrinsic value = 1.1541 - 1.1327 = $0.0214 per euro

Intrinsic value per contract = 0.0214 * 15,000 = $321

This is the minimum value of a put option.

(Note: If the spot price > strike price, then the intrinsic value = 0)

Can you please upvote? Thank You :-)